### v2

@@ 2 Model and Preliminaries
For ease of exposition, we present the classical network revenue management model [^2] There is a special case where the decision-maker only has one type of resource, and the special model is called as single-leg revenue management.. In the main text, we will operate under the assumption that all algorithms are applied to this classical model.
@@
We study two variants of the above model based on the customers’ arrival process. For the first variant (Section 4), we assume that customers arrive sequentially at random; i.e., each customer of type $i$ has an associated arrival rate $\lambda_{i}$. In every round, at most one customer will arrive, with the probability of a type $i$ customer arriving being $\frac{\lambda_{i}}{\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}}$ [^3] Typically, we normalize all arrival rates so that $\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}=1$.. Here, $\lambda_{0}$ represents the rate for no arrivals. In Sections 4.1, and 4.2, we study the case where the arrival rates $\lambda_{i}$ are known; Appendix 13.1 studies the case where the rates $\lambda_{i}$ are unknown. Under stochastic arrivals, regret is commonly used for measuring the performance of an online algorithm. For the second variant (discussed in Section 5), we study the case where an adversary determines both the quantity and sequence of each type of customer, commonly referred to as the worst-case/adversarial scenario. Under adversarial arrivals, the competitive ratio is the metric to study the performance of an online algorithm.
@@ 2.1 Regret and Competitive Ratio Analysis
Next, we consider the adversarial arrival setting. Let $\mathcal{I}$ represent the set of all potential arrival instances. For any $I\in\mathcal{I}$, let $\text{OPT}(I)$ denote the maximum revenue achievable for arrival instance $I$. We denote $\text{Rev}({\pi}(I))$ as the expected revenue elicited by an online algorithm $\pi$ under the specific arrival instance $I$.
@@
The competitive ratio of an algorithm $\pi$ is defined as $\text{CR}(\pi)=\inf_{I\in\mathcal{I}}\frac{\text{Rev}({\pi}(I))}{\text{OPT}(I)}$.
@@
2.2 Popular Algorithm Structures under Stochastic Arrivals
@@
Note that for $\alpha<\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}})$ Algorithm DLP-PA is unfair to all type $i$ customers. This is because if each type $i$ customer is accepted with a probability of $\frac{x^{*}_{i}}{\lambda_{i}}$, it results in a situation where each type $i$ customer is rejected and the prior customer is accepted with probability $\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}})>\alpha$.
Algorithm Resolving DLP (RDLP-PA). The RDLP-PA algorithm addresses DLP’s limitation of not accounting for the demand’s randomness after $t=0$. At a high level, RDLP-PA solves the DLP and employs its optimal primal solution, denoted as $\mathbf{x}^{*}$, to make probabilistic assignments. However, at time point $t^{*}$, the applicability of the initial optimal solution $\mathbf{x}^{*}$ for making decisions may be compromised due to the random noise of the stochastic arrival process and probabilistic decision behavior. Hence, we re-solve the DLP with the remaining capacity and remaining time horizon, and continue making probabilistic assignments by the new optimal solution $\mathbf{\tilde{x}^{*}}$ until the end. For network RM, Reiman and Wang (2008) show that the regret incurred by RDLP-PA is $O(T^{1/3})$. Similarly to the DLP-PA algorithm, the RDLP-PA algorithm is unfair to all type $i$ customers for $\alpha<\min\{\frac{x^{*}_{i}}{\lambda_{i}}(1-\frac{x^{*}_{i}}{\lambda_{i}}),\frac{\tilde{x}^{*}_{i}}{\lambda_{i}}(1-\frac{\tilde{x}^{*}_{i}}{\lambda_{i}})\}$.
@@
We note that, once the bid price for a type $i$ is established, every customer of type $i$ is either accepted (if there is sufficient capacity) or rejected. Therefore, if resource capacity exceeds demand, Algorithm S-BPC is inherently fair because it either accepts or rejects all customers of type $i$. However, if the demand exceeds the resource capacity, Algorithm S-BPC’s first-come-first-serve nature leads to unequal treatment between earlier and later customers.
2.3 Popular Algorithm Structures under Adversarial Arrivals
Booking Limit (BL). In the BL algorithm (Williamson 1992), the supplier assigns a fixed quota/booking limit $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ for each customer type. In the asymptotic setting, all booking limits scale with the resource quantity; hence, $b_{i}=\Theta(m)$ for $i\in[n]$. The algorithm accepts any type $i$ customers provided that there is sufficient resource capacity and the number of accepted type $i$ customers has not yet reached $b_{i}$. Beyond this point, all type $i$ customers are rejected. The BL algorithm is not individually fair; for any type $i$, the $b_{i}^{\text{th}}$ type $i$ customer is accepted while the $(b_{i}+1)^{\text{th}}$ type $i$ customer is rejected.
@@ 2.3 Popular Algorithm Structures under Adversarial Arrivals
The NESTING framework shares some similarities with the BL algorithm. Initially, $n$ quotas $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ need to be established. Following this, $b_{i}$ is assigned as a quota for the total number of type $i$, $i+1$, $\ldots$, $n$ customers. Much like the BL, the NESTING algorithm also lacks individual fairness owing to its quota-based FCFS structure.
@@ 3 The “Grace Period”
(Decreasing Grace Period) The decreasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any $i(k)$ customer that arrives in between round $t_{1}(i)$ and $t_{2}(i)$, $i(k)$ is accepted with a probability of $1-\alpha$ if customer $i(k-1)$ was accepted, and $i(k)$ is rejected if $i(k-1)$ was rejected.
@@ 3.1 Decreasing Grace Period in Network Revenue Management under FCFS
To achieve $(\alpha,\delta)$-fairness in network revenue management under FCFS, we can utilize the decreasing grace period as follows: for any $\alpha,\delta$, let $\gamma=\log_{1-\alpha}\delta$, and assign a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer within the range of $[n]$. Here, $m_{j}(t)$ denotes the remaining capacity of resource $j$ at time $t$. In other words, in order to address the unfairness arising from resource scarcity, it is sufficient for FCFS to have a common decreasing grace period for all customer types. As a result, for FCFS we will drop the dependence on the customer type and denote the grace period as $[t_{1},t_{2}]$. The following theorem states that, by implementing the decreasing grace period as above, we satisfy $(\alpha,\delta)$-fairness, and the revenue loss in comparison to the FCFS algorithm is bounded.
@@
Assigning a decreasing grace period $[t_{1},t_{2}]$, where $t_{1}=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}=T$ to each type $i$ customer, where $\gamma=\log_{1-\alpha}\delta$, FCFS is $(\alpha,\delta)$-fair for network revenue management. Moreover, compared to the FCFS algorithm without grace period, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$, where $\bar{r}=\sup_{i\in[n]}r_{i}$, $\bar{a}=\sup_{i,j}A_{ij}$, and $\underline{a}=\inf_{i,j}A_{ij}$.
To prove Theorem 3.2, we first introduce a lemma below (see Appendix 11 for proof).
Lemma 3.3
Given an online algorithm $\mathcal{A}$, for any customer type $i\in[n]$, and for any $u$, if $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$, then $\mathcal{A}$ satisfies individual fairness (Definition 1.1); namely, for any $u,v\in[n_{i}]$: $\mathbb{P}\left(i(v)\succ_{\mathcal{A}}i(u)\right)\leq\alpha|u-v|$, with probability at least $1-\delta$. Here, the probability mentioned encompasses the overall environmental constraints, including capacity limits, rather than just the algorithm’s decision to accept or reject.
At a high level, the lemma states that for any algorithm $\mathcal{A}$, if for a customer type $i$ the probability that the $u$-th customer gets rejected while the $(u+1)$-th is accepted is upper bounded by $\alpha$ and the same thing holds for customers $u$ and $u-1$, then the algorithm is $(\alpha,\delta)$-fair.
We highlight the intuition of why the revised FCFS is $(\alpha,\delta)$-fair. In the revised algorithm, every customer arriving during $[1,t_{1}-1]$ is accepted, but starting from the period $t_{1}$, the acceptance probability for each customer type progressively decreases. This decrease is conditioned on the preceding customer’s realization. This implies that if the resources are not depleted, the probability that customers of the same type that arrive one after the other get different treatment is bounded by $\alpha$. At round $t_{1}$, the remaining capacity of each resource $j\in[L]$ is at least $\bar{a}n\gamma$. Given that the stopping time (i.e., the time that the algorithm stops accepting them) for each customer type is a geometric random variable, we can apply standard concentration bounds to demonstrate that, with a high probability, all geometric random variables will stop before $\bar{a}n\gamma$ resource $j$ are exhausted for each $j\in[L]$. This concludes the proof that the revised algorithm is $(\alpha,\delta)$-fair.
To obtain a better understanding about the various parameters describing the revenue loss and the efficiency of our proposed algorithm, note that if $\delta=1/m$, and given that $\underline{a}$, $\bar{a}$, $n$ and $\bar{r}$ are constants, Theorem 3.2 implies that the loss in revenue relative to the FCFS algorithm is bounded by $O(\log m)$. Next, we provide the full proof of Theorem 3.2.
Proof 3.4
Proof of Theorem 3.2 First, we show that by assigning a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer, Algorithm FCFS is $(\alpha,\delta)$-fair. By Lemma 3.3, we only need to show that for any customer type $i$, with probability at least $1-\delta$, for any $u\in[n_{i}]$, $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$.
@@
Define event $E$ as $E:=\{\text{no resource is depleted in the time interval }[0,T]\}$. We show that $E$ happens with probability at least $1-\delta$ by showing that the complement of $E$ happens with probability at most $\delta$.
@@
$=(1-\alpha)^{\gamma}$ (4)
$=(1-\alpha)^{\log_{1-\alpha}\delta}=\delta,$
where (3) is due to the pigeonhole principle, and (4) is because for each type $i$, the number of type $i$ customers accepted after the grace period starts is a geometric random variable with success probability $\alpha$. Therefore, the cdf is $1-(1-\alpha)^{\gamma}$.
Conditional on event $E$ happening, for any customer $i(u)$ arriving within $[1,t_{1}]$, the algorithm accepts $i(u)$ with probability $1$. For any customer $i(u)$ arriving within $[t_{1},t_{2}]$, by the definition of the grace period, we obtain that $\mathbb{P}\left(i(u+1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$ and $\mathbb{P}\left(i(u-1)\succ_{\mathcal{A}}i(u)\right)\leq\alpha$.
To complete the proof of the theorem, we need to show that the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$. In the worst case for the revenue, the algorithm rejects all customers after $t_{1}$. By the definition of $t_{1}$, the remaining units for each $j$ are $\bar{a}n\gamma$. Since $\bar{a}n\gamma$ units of resource can serve at most $\frac{\bar{a}}{\underline{a}}n\gamma$ customers, we have that the revenue loss is at most $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$.

### v3

@@ 2 Classical Network Revenue Management Model & Preliminaries
For ease of exposition, we present the classical network revenue management model [^1] There is also a special case where the decision-maker only has one type of resource, and the special model is called as single-leg revenue management.. In the main text, we will operate under the assumption that all algorithms are applied to this classical model. Next, we review the classical network revenue management model.
@@
We study two variants of the above model based on the customers’ arrival process. For the first variant (Section 5), we assume that customers arrive sequentially at random; i.e., each customer of type $i$ has an associated arrival rate $\lambda_{i}$. In every round, at most one customer will arrive, with the probability of a type $i$ customer arriving being $\frac{\lambda_{i}}{\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}}$ [^2] Typically, we normalize all arrival rates so that $\sum_{i^{\prime}=0}^{n}\lambda_{i^{\prime}}=1$.. Here, $\lambda_{0}$ represents the rate for no arrivals. In Sections 5.1, and 5.2, we study the case where the arrival rates $\lambda_{i}$ are known; Section 5.3 studies the case where the rates $\lambda_{i}$ are unknown. Under stochastic arrivals, regret is commonly used for measuring the performance of an online algorithm. For the second variant (discussed in Section 6), we study the case where an adversary determines both the quantity and sequence of each type of customer, commonly referred to as the worst-case/adversarial scenario. Under adversarial arrivals, the competitive ratio is the metric to study the performance of an online algorithm.
@@ 2.1 Regret and Competitive Ratio Analysis
Next, we consider the adversarial arrival setting. Let $\mathcal{I}$ represent the set of all potential arrival instances. For any $I\in\mathcal{I}$, let $\text{Rev}^{\star}(I)$ denote the maximum revenue achievable for arrival instance $I$. We denote $\text{Rev}({\pi}(I))$ as the expected revenue elicited by an online algorithm $\pi$ under the specific arrival instance $I$.
@@
The competitive ratio of an algorithm $\pi$ is defined as $\text{CR}(\pi)=\inf_{I\in\mathcal{I}}\frac{\text{Rev}({\pi}(I))}{\text{Rev}^{\star}(I)}$.
@@
2.2 Review: Popular Algorithm Structures under Stochastic Arrivals
In this section, we provide a review of existing algorithmic frameworks designed for stochastic arrival processes. These algorithms can be broadly categorized into two main approaches:
  1. 1.
    Primal Methods (Section 2.2.1): These algorithms rely on the primal solution of the linear programming formulation to guide decision-making.
  2. 2.
    Dual Methods (Section 2.2.2): These approaches leverage the dual solution of the linear programming formulation to inform allocation decisions.
2.2.1 Primal Methods
@@
Algorithm Resolving DLP (RDLP-PA). The RDLP-PA algorithm addresses DLP’s limitation of not accounting for the demand’s randomness after $t=0$. At a high level, RDLP-PA solves the DLP and employs its optimal primal solution, denoted as $\mathbf{x}^{*}$, to make probabilistic assignments. However, at time point $t^{*}$, the applicability of the initial optimal solution $\mathbf{x}^{*}$ for making decisions may be compromised due to the random noise of the stochastic arrival process and probabilistic decision behavior. Hence, we re-solve the DLP with the remaining capacity and remaining time horizon, and continue making probabilistic assignments by the new optimal solution $\mathbf{\tilde{x}^{*}}$ until the end. Reiman and Wang (2008), Jasin and Kumar (2012), Bumpensanti and Wang (2020) apply this approach with different resolving times in revenue management and achieve $o(\sqrt{T})$ regret.
An alternative version of RDLP-PA, proposed by Vera and Banerjee (2019), takes a different approach to make online decisions after resolving the DLP at every single time period. Instead of using probabilistic assignments, this version employs the compensated coupling technique, making deterministic accept/reject decisions based on the resolved optimal primal solutions. This method achieves a regret bound comparable to the standard RDLP-PA. We will discuss in Section 3.1 its inability to alleviate our fairness concerns.
2.2.2 Dual Methods
Primal methods in revenue management use the dual solution of a linear program to guide allocation decisions.
@@ 2.2.2 Dual Methods
Algorithm Dynamic Bid Price Control (BPC-OGD). Similarly to primal methods, dual-based approaches can also dynamically update the dual variables at each time period. However, unlike primal methods, which often require resolving an LP, dual methods are well-known for their computational efficiency. Rather than recomputing the optimal dual solution at every step, these methods rely on first-order online optimization techniques to update the bid price dynamically. For example, Li et al. (2020), Sun et al. (2020), Balseiro et al. (2023b) propose a dynamic bid price control approach, leveraging online gradient descent or online mirror descent (BPC-OGD). By continuously adjusting the bid prices, these algorithms set a dynamic threshold price for each resource, making deterministic accept/reject decisions for every customer.
Compared to S-BPC, BPC-OGD achieves the same $O(\sqrt{T})$ regret bound under stochastic arrivals while offering two key advantages: (i) No prior knowledge of arrival rates: the algorithm remains effective even when the arrival rates $\lambda_{i}$ are entirely unknown. (ii) No need to solve any primal or dual LP. However, as we discuss in Section 3.1, BPC-OGD does not satisfy our fairness desiderata.
2.3 Review: Popular Algorithm Structures under Adversarial Arrivals
Booking Limit (BL). In the BL algorithm (Williamson 1992), the supplier assigns a fixed quota/booking limit $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ for each customer type. In the asymptotic setting, all booking limits scale with the resource quantity; hence, $b_{i}=\Theta(m)$ for $i\in[n]$. The algorithm accepts any type $i$ customers provided that there is sufficient resource capacity and the number of accepted type $i$ customers has not yet reached $b_{i}$. Beyond this point, all type $i$ customers are rejected.
@@ 2.3 Review: Popular Algorithm Structures under Adversarial Arrivals
The NESTING framework shares some similarities with the BL algorithm. Initially, $n$ quotas $\mathbf{b}=\{b_{1},b_{2},\ldots,b_{n}\}$ need to be established. Following this, $b_{i}$ is assigned as a quota for the total number of type $i$, $i+1$, $\ldots$, $n$ customers.
@@
3 Definition of Individual Fairness Metrics
In Section 1, we conceptually introduced individual fairness in quantity-based revenue management from the customer’s perspective. Recall, a customer does not perceive unfairness simply from being denied a resource; rather, unfairness arises when some customers receive the resource while others, requesting the same resource under similar conditions, do not. To quantify this disparity in treatment, we use probability as a measure of fairness. Ensuring individual fairness requires minimizing the probability that two similar customers receive different treatment. Having established the classical quantity-based network revenue management model in Section 2, we now formally define the individual fairness metric.
Definition 3.1 (Individual Fairness in Revenue Management)
Let $n$ be the number of customer types, with each type $i\in[n]$ having $n_{i}$ customers arriving. Let $[n_{i}]$ be the set which contains all type $i$ customers. An algorithm $\mathcal{A}$ is $(\alpha,\delta)$-fair for any choice of $\alpha,\delta\in(0,1)$ if for each customer of type $i$, with a probability at least $1-\delta$, for any $u,v\in[n_{i}]$,
$\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]\leq\alpha\cdot d(u,v),$ (3)
where $i(u),i(v)$ denote the $u$-th and $v$-th customer’s of type $i$, and $i(v)\succ_{\mathcal{A}}i(u)$ indicates that algorithm $\mathcal{A}$ does not accept $i(u)$ given that it accepts customer $i(v)$, namely, $\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]=\mathbb{P}\left[i(v)\text{ is accepted}|i(u)\text{ is rejected}\right]$. $d(u,v)$ is a general time-distance function that satisfies the following properties:
  1. 1.
    $d(u,v)$ is increasing with respect to $|u-v|$.
  2. 2.
    $d(u,v)=1$ if $|u-v|=1$.
  3. 3.
    For any $\alpha\in(0,1)$ chosen by the seller, there exist $u,v$ within the seller-buyer interaction horizon $T$ such that $d(u,v)=1/\alpha$. Specifically, let $w_{\alpha}>1$ (independent of $T$) be the smallest positive integer such that $|u-v|=w_{\alpha}$ and $d(u,v)\geq 1/\alpha$,
This definition ensures that fairness constraints hold with high probability, allowing for more flexible and practical revenue management policies while mitigating inefficiencies caused by extreme arrival sequences. At a high level, property 3 implies that customers do not really care about being treated differently with customers that are very far into the future. To better understand the implications and design choices behind this fairness metric, we include an extensive discussion in Section 10.
3.1 Unfairness in Classical RM Algorithms
After introducing the fairness metric in Definition 3.1, we show that all classical revenue management (RM) algorithm structures can violate individual fairness for certain parameter choices of $(\alpha,\delta)$. We classify the sources of unfairness into three categories:
(i) Unfairness due to resource capacity limits. This form of unfairness arises from the fundamental constraint that resources are limited. It affects all algorithms: for example, consider two type $i$ customers who arrive close to each other. If the earlier customer arrives when the resource is already depleted, they will be rejected, while the later customer may have been accepted had they arrived earlier. Even though these customers arrive close in time and request the same resource, one may be accepted while the other is inevitably rejected simply because the resource becomes unavailable in between their arrivals.
(ii) Unfairness due to randomized individual decisions. Algorithms such as DLP-PA and RDLP-PA fall into this category. These algorithms assign each type $i$ customer an acceptance probability of $\frac{x_{i}^{*}}{\lambda_{i}}$, independently of others. While this appears fair from the decision-maker’s perspective, it results in observable disparities. For instance, two consecutive type $i$ customers may receive different outcomes—one accepted, the other rejected—with a non-negligible probability. If $\alpha$ is smaller than this disparity probability ($1-\frac{x_{i}^{*}}{\lambda_{i}}$), the algorithm violates the fairness constraint.
(iii) Unfairness due to deterministic individual decisions. This issue arises in algorithms that make accept/reject decisions for each individual based on deterministic rules. Notable examples include the alternative version of RDLP-PA proposed in Vera and Banerjee (2019) and the BPC-OGD algorithm. These algorithms resolve a linear program (either primal or dual) at each time step and deterministically decide whether to accept or reject a customer. As a result, consecutive customers of the same type may experience alternating treatment patterns—e.g., one accepted, the next rejected—even when they arrive in close succession. This deterministic oscillation leads to a violation of the fairness metric for any $\alpha<1$.
3.2 Objective: Enhancing Classical RM Algorithms to Satisfy Fairness Constraints with Minimal Efficiency Loss
Having introduced the fairness metric in Definition 3.1, we now formalize the main goal of this paper: to enhance classical revenue management (RM) algorithms so that they satisfy the fairness constraint without significantly sacrificing asymptotic performance.
Specifically, we aim to enhance the algorithmic structures discussed in Section 2—including primal-based, dual-based, and booking-limit-based methods. Roughly speaking, we define “enhancement” as the ability to maintain the same computational complexity as the original algorithm while also satisfying our fairness definition. We aim for our enhanced algorithms to maintain near-optimal regret for stochastic settings, and a similar-to-the-original competitive ratio.
@@ 4 The “Grace Period”
(Decreasing Grace Period) The decreasing grace period $[t_{1}(i),t_{2}(i)]$ for type $i$ customers is defined as: for any $i(k)$ customer that arrives between round $t_{1}(i)$ and $t_{2}(i)$, $i(k)$ is accepted with a probability of $1-\beta$ if customer $i(k-1)$ was accepted, and $i(k)$ is rejected if $i(k-1)$ was rejected, where $\beta=1-(1-\alpha)^{1/w_{\alpha}}$ and $w_{\alpha}$ is defined in the third property of $d(u,v)$ in Definition 3.1.333Note that $w_{\alpha}$ is a parameter of the problem controlled by the seller. Specifically, the seller originally chooses the target $\alpha$ together with the distance metric $d(u,v)$ and then just solves the equation $d(u,v)=1/\alpha$ to identify $w_{\alpha}$. In the decreasing grace period design, each customer’s acceptance depends only on the outcome of the immediately preceding customer, and is otherwise probabilistically independent.
@@ 4.1 Decreasing Grace Period in Network Revenue Management under FCFS
To achieve $(\alpha,\delta)$-fairness in network revenue management under FCFS, we can utilize the decreasing grace period as follows: for any $\alpha,\delta$, let $\gamma=\log_{1-\beta}\delta$, and assign a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer within the range of $[n]$. Here, $m_{j}(t)$ denotes the remaining capacity of resource $j$ at time $t$. In other words, in order to address the unfairness arising from resource scarcity, it is sufficient for FCFS to have a common decreasing grace period for all customer types. As a result, for FCFS we will drop the dependence on the customer type and denote the grace period as $[t_{1},t_{2}]$. The following theorem states that, by implementing the decreasing grace period as above, we satisfy $(\alpha,\delta)$-fairness, and the revenue loss in comparison to the FCFS algorithm is bounded.
@@
Assigning a decreasing grace period $[t_{1},t_{2}]$, where $t_{1}=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}=T$ to each type $i$ customer, where $\gamma=\log_{1-\beta}\delta$, $\beta=1-(1-\alpha)^{1/w_{\alpha}}$ and $w_{\alpha}$ is defined in the third property of $d(u,v)$ in Definition 3.1. FCFS is $(\alpha,\delta)$-fair for network revenue management. Moreover, compared to the FCFS algorithm without grace period, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{r}$, where $\bar{r}=\sup_{i\in[n]}r_{i}$, $\bar{a}=\sup_{i,j}A_{ij}$, and $\underline{a}=\inf_{i,j}A_{ij}$.
Proof 4.3
Proof of Theorem 4.2 We first show that without the resource capacity constraint, for any $u,v\in[n_{i}]$, the decreasing grace period can guarantee that
$\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]\leq\alpha\cdot d(u,v).$ (4)
W.L.O.G, we assume that $u>v$ and denote $x=u-v$. If $x\geq w_{\alpha}$, Inequality (4) holds since $\alpha d(u,v)\geq 1$; this is because $d(u,v)$ is increasing in $|u-v|$ and $d(u,v)=1/\alpha$ when $|u-v|=w_{\alpha}$. If $x<w_{\alpha}$, recall the definition of the decreasing grace period: if the previous customer is accepted, the next is accepted with probability $1-\beta$; otherwise, the next is rejected. Thus, if customer $i(v)$ is accepted and $i(u)$ is rejected, it must be that some customer in $\{i(v+1),\ldots,i(u)\}$ is the first to be rejected after a sequence of accepted customers. Specifically, for each $k\in\{v+1,\ldots,u\}$, we consider the case where $i(v+1),\ldots,i(k-1)$ are accepted and $i(k)$ is rejected. We then sum the probabilities of all such scenarios:
$\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]$ $=\mathbb{P}\left[\text{one of the customers in }\{i(v+1),\ldots,i(u)\}\text{ is the first to be rejected}|i(v)\text{ is accepted}\right]$
$\leq\sum_{j=0}^{x-1}\mathbb{P}\left[i(v+1),i(v+2),\ldots,i(v+j)\text{ is accepted and }i(v+j+1)\text{ is rejected}|i(v)\text{ is accepted}\right]$
$=\sum_{j=0}^{x-1}\prod_{\ell=1}^{j}\mathbb{P}\left[i(v+\ell)\text{ is accepted}\,\middle|\,i(v),i(v+1),\ldots,i(v+\ell-1)\text{ is accepted}\right]$
$\quad\quad\cdot\mathbb{P}\left[i(v+j+1)\text{ is rejected}\,\middle|\,i(v),i(v+1),\ldots,i(v+j)\text{ is accepted}\right]$
$=\sum_{j=0}^{x-1}\prod_{\ell=1}^{j}\mathbb{P}\left[i(v+\ell)\text{ is accepted}\,\middle|\,i(v+\ell-1)\text{ is accepted}\right]$
$\quad\quad\cdot\mathbb{P}\left[i(v+j+1)\text{ is rejected}\,\middle|\,i(v+j)\text{ is accepted}\right]$
$=\sum_{j=0}^{x-1}(1-\beta)^{j}\beta$
$=\beta\frac{1-(1-\beta)^{x}}{1-(1-\beta)}$
$=1-(1-\beta)^{x}.$
Since $1-(1-\beta)^{x}$ is increasing in $x$, we have for $x<w_{\alpha}$,
$1-(1-\beta)^{x}<1-(1-\beta)^{w_{\alpha}}=1-\left(1-\left(1-(1-\alpha)^{1/w_{\alpha}}\right)\right)^{w_{\alpha}}=1-(1-\alpha)=\alpha.$ (5)
As $\alpha d(u,v)$ is also an increasing function with $x=u-v$ and $\alpha d(u,v)=\alpha$ when $x=1$, by Equation (5), we have
$1-(1-\beta)^{x}<\alpha\leq\alpha d(u,v).$
Therefore, we have for any $u,v\in[n_{i}]$, the decreasing grace period satisfies Equation (4). If $u<v$, a symmetric argument also holds and we omit the proof.
@@
Define event $E$ as $E:=\{\text{no resource is depleted in the time interval }[0,T]\}$. The above analysis shows that if $E$ happens, then the fairness constraint holds. Next, we show that $E$ happens with probability at least $1-\delta$ by showing that the complement of $E$ happens with probability at most $\delta$.
@@
$=(1-\beta)^{\gamma}$ (7)
$=(1-\beta)^{\log_{1-\beta}\delta}=\delta,$
where (6) is due to the pigeonhole principle, and (7) is because for each type $i$, the number of type $i$ customers accepted after the grace period starts is a geometric random variable with success probability $\beta$. Therefore, the cdf is $1-(1-\beta)^{\gamma}$.
[…]
We highlight the intuition of why the revised FCFS is $(\alpha,\delta)$-fair. In the revised algorithm, every customer arriving during $[1,t_{1}-1]$ is accepted, but starting from the period $t_{1}$, the acceptance probability for each customer type progressively decreases. This decrease is conditioned on the preceding customer’s realization. This implies that if the resources are not depleted, the probability that customers of the same type that arrive one after the other get different treatment is bounded by $\beta=1-(1-\alpha)^{1/w_{\alpha}}$. The choice of $\beta$ ensures that, for any pair of customers $i(u)$ and $i(u+x)$ with $x\in[w_{\alpha}]$, the conditional probability that $i(u)$ is accepted and $i(u+x)$ is rejected is at most $\alpha$, thereby satisfying the fairness constraint. For $x>w_{\alpha}$, by the first and third properties of $d(u,v)$ in Definition 3.1, we have $\alpha\cdot d(u,u+x)\geq 1$, so the fairness constraint is automatically satisfied since the probability is always upper bounded by $1$. At round $t_{1}$, the remaining capacity of each resource $j\in[L]$ is at least $\bar{a}n\gamma$. Given that the stopping time (i.e., the time that the algorithm stops accepting them) for each customer type is a geometric random variable, we can apply standard concentration bounds to demonstrate that, with a high probability, all geometric random variables will stop before $\bar{a}n\gamma$ resource $j$ are exhausted for each $j\in[L]$. This concludes the proof that the revised algorithm is $(\alpha,\delta)$-fair.
To obtain a better understanding about the various parameters describing the revenue loss and the efficiency of our proposed algorithm, note that if $\delta=1/m$, and given that $\underline{a}$, $\bar{a}$, $n$ and $\bar{r}$ are constants, Theorem 4.2 implies that the loss in revenue relative to the FCFS algorithm is bounded by $O(\log m)$. The full proof of Theorem 4.2 can be found in Appendix 14.